a youthful person run at 7km/h in a north- west direction across the derk of a ship in which is streaming due east at 40 km/h .ifind the velocity of the boy relative to the Sea
ii,And the velocity of the sea relative to the boy
1
Expert's answer
2021-07-19T09:48:51-0400

i)


"v=\sqrt{(7\cos 45)^2+(40-7\sin 45)^2}=35.4\frac{km}{h}\\
\theta=\arctan{\frac{7\cos 45}{(40-7\sin 45)}}=8.0\degree\text{ N of E}"

ii)


"v=35.4\frac{km}{h}\\
\theta=8.0\degree\text{ S of W}"